The area of a circle is directly proportional to the square of its radius. The area of a circle whose diameter is 40 cm is 1256 sq cm. What is the area of a circle whose diameter is 20 cm? 10 cm? 12 cm? 30 cm?
Answer by Cromlix(4381)

Hi there,
The area of a circle is directly
proportional to the square of its radius.
A = kr^2
Where k = constant of proportionality
A = kr^2
1256 = k20^2
k = 1256/400
k = 3.14
Formula:
A = 3.14r^2
Problem:
A = 3.14 x 10^2
A = 3.14 x 100
A = 314 sq cm.
Hope this helps :-)